NEXT
Make New Entry,
Make Followup Entry
User name riordan
Log entry time 21:21:53 on March 10, 2009
Entry number 265562
keyword=Summary for March 9 Swing - March 10 Day
Summary for Mar 9 swing - Mar 10 Day
MCC continues to try and get beam restored
Installed material around Compton
Waiting for beamline vacuum to be restored
Beam in Use
Shift |
ABU |
BANU |
BNA |
2009/03/09 Swing | 0.00 |
0.00 | 8.00 |
2009/03/10 Owl | 0.00 |
0.00 | 8.00 |
2009/03/10 Day | 0.00 |
0.00 | 8.00 |
TOTAL | 0.00 |
0.00 | 24.00 |
Main events/issues:
- Mar 9
- Beam still down due to 1L04
- Mar 10
- Installed additional material around the Compton electron
detector. Waiting to restore vacuum.
Plans for the next few days:
- Positive polarity running on BigBite and LHRS when beam is restored
- Continue 4-pass transverse production for as long as possible
A copy of this log entry has been emailed to: brads, jpchen, jiang, meziani, choi, riordan